In the YDSE the slits are 0.5mm apart and inteference is observed on a screen placed at a distance of 100cm from the slits . it is found that the 9th bright fringe is at a distane of 9mm from the second dark fringe from the centre of the fringe pattern. what is the wavelnght of light used ?

Grade:12

1 Answers

Arun
6 years ago
d = 0.5 mm
= 0.5 × 10–3 m
= 5 × 10–4 m
D = 100 cm = 1 m
 X9
 – X′2
= 8.835 mm
[Assuming both bright and dark fringes
are on same side of centre of fringe
pattern]
= 8.835 × 10–3 m
To Find :
λ = ?
Formula :
λ =
Xd/D
Since distance of nth bright band
Xn = nλD/d
X9 = 9λD/d
X2' = (2*2- 1) λD/2d
X2' = 3λD/2d
X9 - X2' = 15λD/2d
Now you can find λ
Where X9-X2' = 9 mm
Hope it helps